# Learning to Plan with Personalized Preferences

### v2

@@ 1 Introduction
The field of embodied Artificial Intelligence (AI) is rapidly advancing, driven by significant progress in foundation models for vision and language (Bommasani et al.,, 2021; Peng et al.,, 2023; Achiam et al.,, 2023; Bai et al.,, 2023). These advances enable AI systems to autonomously collaborate with or assist humans in daily tasks, particularly in domestic settings (Driess et al.,, 2023; Leal et al.,, 2023; Zitkovich et al.,, 2023; Ahn et al.,, 2024). However, recent approaches utilizing natural language instructions (Mu et al.,, 2023; Zitkovich et al.,, 2023; Singh et al.,, 2023) face fundamental limitations in capturing human preferences (Zhu et al.,, 2016). While natural language is our primary means of communication, its inherent ambiguity creates a gap between instructions and intended executions (Yuan et al.,, 2022; Jiang et al.,, 2022; 2021; Yuan et al.,, 2020). For instance, when a user requests help in preparing an apple, the agent needs to understand specific preferences about apple selection, washing requirements, cutting style, and container choice—details that vary significantly across individuals; see also Figure 1 for a graphical illustration.
@@
Learning human preferences in real-world settings presents unique challenges (Peng et al.,, 2024). Humans typically communicate their needs succinctly (Levinson,, 1983), without exhaustive preference details (Lichtenstein and Slovic,, 2006), and many preferences include unconscious or instinctive elements difficult to articulate (Epstein,, 1994; Simonson,, 2008). A more practical approach is to infer preferences from observed human choices and decision-making patterns, as illustrated in Figure 1, where a robot assistant can learn users’ preferences and behavior habits from previous observations.
@@ 4.1 Definition of Preferences
We organize preferences in a three-tiered hierarchical structure that captures varying degrees of specificity across tasks. Figure 2 provides an overview of all preferences and their distribution, while Figure 3 illustrates concrete examples of preferences and corresponding agent actions. The 290 preferences are distributed across three levels: 80 sequence-level, 135 option-level, and 75 action-level preferences.
@@ 4.2 Constructing PbP Test Set
Our PbP benchmark includes a default test set for systematic model evaluation. Following the formulation in Section 3, we structure PbP tasks as few-shot learning-from-demonstration problems. Each test point comprises several (typically three) unique demonstrations with egocentric observations of action sequences and their corresponding preference labels. As illustrated in Figure 4, a demonstration includes an egocentric video of agent activity, a bird’s-eye-view map tracking agent position, and frame-level action annotations. We also provide third-person view recordings for enhanced visualization. We prioritize the egocentric perspective for two reasons: 1) it offers a clear view with minimal occlusions, and 2) it aligns with human perception, facilitating transfer to real-world data from head-mounted devices.
@@ 4.3 Models
To assess pure language understanding, we evaluate advanced LLMs including Llama3 (Touvron et al.,, 2023) and GPT-4 (Achiam et al.,, 2023) using only action sequence inputs. This approach treats actions as high-level abstractions of egocentric videos, reducing visual complexity while maintaining task semantics. We benchmark Llama3-8B as our baseline against GPT-4-Turbo as the current state-of-the-art, employing prompt designs informed by the OpenAI Cookbook for optimal few-shot performance. Detailed model architectures and prompt engineering strategies are discussed in Appendix D.
@@ 5.2 End-to-end Action Preference Learning
As shown in Table 1 (the End-to-end row), video-based models produce Levenshtein distances approaching the average ground truth sequence lengths (15.80 at option level, 35.87 at sequence level). These high distances indicate that the models generate predominantly inconsistent action sequences, suggesting a failure to grasp preferences embedded in demonstration videos. While symbol-based models show modest improvements, their performance gains remain limited.
Table 1: Levenshtein distance between generated and ground truth action sequences. End-to-end represents models directly generating action sequences from demonstration-preference pair examples. Second-stage indicates generation using both demonstrations and previously inferred preference labels based on demonstrations. Second-stage (gt) uses demonstrations alongside ground truth preference labels for sequence generation.
@@
ViViT LLaVA-Next EILEV GPT-4V Llama3-8B GPT-4
Option Level End-to-end 15.49$\pm$1.29 15.94$\pm$3.41 12.88$\pm$2.20 15.63$\pm$2.31 14.74$\pm$3.21 12.23$\pm$2.96
Second-stage - 12.46$\pm$3.23 12.89$\pm$3.74 8.37$\pm$2.19 9.67$\pm$5.16 2.23$\pm$2.79
Second-stage (gt) - 3.28$\pm$5.29 11.18$\pm$4.20 1.26$\pm$2.55 8.22$\pm$5.58 0.12$\pm$3.12
Sequence Level End-to-end 34.04$\pm$11.84 34.76$\pm$11.25 33.10$\pm$12.21 33.75$\pm$11.15 31.79$\pm$7.32 27.85$\pm$6.57
Second-stage - 30.02$\pm$13.54 33.03$\pm$13.61 27.52$\pm$9.48 25.46$\pm$5.93 16.45$\pm$4.00
Second-stage (gt) - 18.92$\pm$14.18 26.57$\pm$12.21 11.36$\pm$8.05 19.02$\pm$7.10 12.29$\pm$3.12
Overall End-to-end 24.76 25.35 22.99 24.69 23.26 20.04
Second-stage - 21.24 22.96 17.94 17.56 9.34
Second-stage (gt) - 11.10 18.88 6.31 13.62 6.21
@@ 5.3 Two-Stage Learning-Planning
Given the limitations of end-to-end learning, we implement a two-stage approach to decompose the preference learning problem. The first stage focuses on preference prediction, where we provide models with auxiliary preference token labels and train them to predict hidden preferences explicitly. These preference tokens, as discussed in Section 4.1, maintain sufficient semantic content for translation into primitive actions.
Results from the first stage (Table 2) reveal significant performance variations across models. At the option level, GPT-4V achieves superior performance with 48.48% accuracy, demonstrating strong capability in interpreting demonstrated preferences. Among symbol-based models, the stark contrast between DAG-Opt’s limited performance and the improved results from Llama3-8B and GPT-4 highlights the advantage of next-token prediction over dependency learning for preference inference. Models with language components consistently show improved preference understanding compared to end-to-end learning.
Table 2: Preference prediction accuracy in few-shot setting.
@@
ViViT LLaVA-Next EILEV GPT-4V DAG-Opt Llama3-8B GPT-4
Option Level 9.38 36.87 38.33 48.48 10.15 72.98 86.27
Sequence Level 4.24 24.85 32.69 37.50 13.49 67.18 68.42
Overall 6.81 30.86 35.51 42.99 11.82 70.08 77.34
The second stage involves generating action sequences based on both demonstrations and predicted preference labels from the first stage, introducing potential error propagation. Results in Table 1 (Second-stage row) and Figure 5 show significant improvements when models receive explicit preferences. For comprehensive evaluation, we include planning results using ground truth preference labels (Second-stage (gt) row). GPT-4V and GPT-4 achieve near-zero Levenshtein distances, indicating almost perfect alignment with ground truth action sequences.
@@
To isolate the impact of prior knowledge versus in-context learning, we conduct ablation studies by removing demonstrations and testing preference prediction on isolated test sequences. Results in Table 3 show significant performance degradation compared to few-shot learning (Table 2), particularly at the sequence level. This suggests that while models may encode basic task-specific preferences, they rely heavily on demonstrations to recognize complex preference patterns in varied sequences.
@@
Figure 5: Levenshtein distance between generated and ground truth action sequences. Results shown for both (a) option level and (b) sequence level under two conditions: End-to-end bars represent direct sequence generation from previous observations, while Second-stage bars show performance when models receive predicted preference labels. The line indicates average ground truth sequence length (option level: 15.80, sequence level: 35.87). Lower distances indicate better performance. Results demonstrate significantly improved performance under the two-stage approach compared to end-to-end generation.
Table 3: Preference prediction accuracy in ablative setting. Evaluation of models’ performance when demonstrations are removed, testing their ability to predict preferences from test sequences alone.
VIDEO-BASED INPUT SYMBOL-BASED INPUT
ViViT LLaVA-Next EILEV GPT-4V DAG-Opt Llama3-8B GPT-4
Option Level 9.16 15.47 4.77 29.42 3.84 39.50 73.87
Sequence Level 4.38 8.13 0.00 0.00 1.28 6.25 9.42
Overall 6.77 11.8 2.38 14.71 2.56 22.88 41.64
@@ 5.4 Generalization
While human actions may vary across different objects and scenes, underlying preferences often remain consistent. We evaluate the models’ ability to generalize preference learning across varying visual contexts. The original test set inherently tests generalization by randomly sampling scenes and objects when rendering video demonstrations for each preference. To gain additional insights, we conduct complementary experiments with controlled conditions where demonstration and test videos are rendered in identical rooms with the same objects. This controlled setting enables direct performance comparisons under consistent conditions. We evaluate EILEV, LLaVA, and GPT-4 series models on this variant of PbP, as these models previously demonstrated strong few-shot reasoning capabilities. Results are summarized in Table 4.
Table 4: Models’ generalization ability. direct denotes experiments conducted without generalization. orig denotes the original experiments conducted with generalization cases. Also the accuracy of preference prediction.
LLaVA-Next EILEV GPT-4V GPT-4
Option Level direct 33.25 46.93 53.24 86.32
Option Level orig 36.87 38.33 48.48 86.27
Sequence Level direct 33.12 37.53 39.42 70.27
Sequence Level orig 24.85 32.69 37.50 68.42
@@
Symbol-based reasoning (GPT-4) demonstrates consistent performance regardless of scene or object variations, while vision-based models show greater sensitivity to scene changes. This distinction stems from the nature of our predefined preferences, which are sufficiently abstract and general to apply across diverse scenes and objects. Vision-based models, however, tend to anchor their few-shot learned preferences to specific visual features of scenes or objects. When these visual elements change, preference recognition accuracy may deteriorate. This contextual dependence remains a persistent challenge for vision-based models, which often overfit to scene-specific features from training videos.
Analysis of test points across direct and gen conditions (Figure 6) reveals two key findings: (i) Preference learning performance correlates with scene characteristics, with certain scenes proving consistently challenging across both conditions. (ii) While direct cases show better performance overall, failure patterns differ between conditions, particularly for vision-based models. This suggests models rely heavily on visual context consistency–including object arrangement and scene layout–for accurate predictions, indicating potential superficial learning rather than true preference understanding. Symbol-based reasoning maintains robust performance across varied scenes due to the general nature of predefined preferences, whereas vision-based models’ strong dependence on specific visual contexts limits their generalization capability.
@@ 5.5 Ablations on Demonstration Numbers
We examine the effect of demonstration quantity on model performance through an ablation study (Figure 7). Results show that increasing demonstration numbers generally improves preference learning and planning effectiveness. This improvement is most evident in second-stage planning, where models achieve lower sequence distances by more accurately replicating human actions. Models like GPT-4, Llama3, and EILEV show consistent performance gains with additional demonstrations. However, we observe that excessive demonstrations (e.g., 5-demo cases for GPT-4 and EILVE) can sometimes impair first-stage prediction accuracy. Despite these occasional exceptions, the overall trend confirms our intuition: more demonstrations enhance learning and planning performance. These findings highlight the importance of demonstration quantity in developing effective personalized planning systems that align with user preferences.
@@ Appendix A Dataset Card
      Each instance contains an egocentric video of an agent’s activity, its bird’s-eye-view map of the position of the agent, and a frame-level textual annotation of the current action, as shown in Figure 4. Additionally, we provide a rendered third-person view of the entire process.
@@ B.1 Preferenes
See Table A1 for the preference statistics in PbP.
@@ B.2 Actions
See Table A2 for the action statistics in PbP. We implement 17 action primitives in PbP to assist with model planning and dataset rendering. These action primitives have parameters that simplify tasks and are considered the lowest-level actions. Each sub-task contains 8 to 20 such lowest-level actions. Generally, most of these actions consist of two parts: the robot movement part and the arm (gripper) execution part. For robot movement, we use the A* algorithm to find paths and avoid collisions. We build a connection map during scene initialization for navigation, taking the robot’s width into consideration. For the arm (gripper) execution, we primarily use the IK algorithm to compute arm movements. However, since IK cannot handle complex tasks, such as picking objects from the fridge, we also leverage the Open Motion Planning Library (OMPL) planner (Sucan et al.,, 2012) with forward planning to assist in planning the arm positions.
@@ B.3 More Dataset Details and Discussion
The process of producing data is mainly explained in Section 4.2. In summary, we follow the order of “sample preference - sample scene - sample objects to be manipulated - generate actions guided by a rule-based planner”.
@@
Considering that preferences in household activities are not only multi-dimensional but also hierarchical, we first define a hierarchy of preferences from the perspective of how things happen in a life scenario, that is, from each specific action to a sub-task consisting of several actions, and then to the sequence combining these sub-tasks. The next step is to expand each level with typical tasks and actions. The detailed definition of the 290 preferences can be found in Section 4.1.
@@ D.6 Llama3-8B
D.7 GPT-4
We use “gpt-4-turbo-2024-04-09” with a temperature of 0.05. The textual prompts are as follows:

### v3

@@ 1 Introduction
The field of embodied Artificial Intelligence (AI) is rapidly advancing, driven by significant progress in foundation models for vision and language (Bommasani et al.,, 2021; Peng et al.,, 2023; Achiam et al.,, 2023; Bai et al.,, 2023). These advances enable AI systems to autonomously collaborate with or assist humans in daily tasks, particularly in domestic settings (Driess et al.,, 2023; Leal et al.,, 2023; Zitkovich et al.,, 2023; Ahn et al.,, 2024). However, recent approaches utilizing natural language instructions (Mu et al.,, 2023; Zitkovich et al.,, 2023; Singh et al.,, 2023) face fundamental limitations in capturing human preferences (Zhu et al.,, 2016). While natural language is our primary means of communication, its inherent ambiguity creates a gap between instructions and intended executions (Yuan et al.,, 2022; Jiang et al.,, 2022; 2021; Yuan et al.,, 2020). For instance, when a user requests help in preparing an apple, the agent needs to understand specific preferences about apple selection, washing requirements, cutting style, and container choice—details that vary significantly across individuals; see also Figure˜1 for a graphical illustration.
@@
Learning human preferences in real-world settings presents unique challenges (Peng et al.,, 2024). Humans typically communicate their needs succinctly (Levinson,, 1983), without exhaustive preference details (Lichtenstein and Slovic,, 2006), and many preferences include unconscious or instinctive elements difficult to articulate (Epstein,, 1994; Simonson,, 2008). A more practical approach is to infer preferences from observed human choices and decision-making patterns, as illustrated in Figure˜1, where a robot assistant can learn users’ preferences and behavior habits from previous observations.
@@ 4.1 Definition of Preferences
We organize preferences in a three-tiered hierarchical structure that captures varying degrees of specificity across tasks. Figure˜2 provides an overview of all preferences and their distribution, while Figure˜3 illustrates concrete examples of preferences and corresponding agent actions. The 290 preferences are distributed across three levels: 80 sequence-level, 135 option-level, and 75 action-level preferences.
@@ 4.2 Constructing PbP Test Set
Our PbP benchmark includes a default test set for systematic model evaluation. Following the formulation in Section˜3, we structure PbP tasks as few-shot learning-from-demonstration problems. Each test point comprises several (typically three) unique demonstrations with egocentric observations of action sequences and their corresponding preference labels. As illustrated in Figure˜4, a demonstration includes an egocentric video of agent activity, a bird’s-eye-view map tracking agent position, and frame-level action annotations. We also provide third-person view recordings for enhanced visualization. We prioritize the egocentric perspective for two reasons: 1) it offers a clear view with minimal occlusions, and 2) it aligns with human perception, facilitating transfer to real-world data from head-mounted devices.
@@ 4.3 Models
To assess pure language understanding, we evaluate advanced LLMs like Llama3 (Touvron et al.,, 2023) and GPT-4.1 (Achiam et al.,, 2023) using only action sequence inputs. This approach treats actions as high-level abstractions of egocentric videos, reducing visual complexity while maintaining task semantics. We benchmark Llama3-8B as our baseline against GPT-4.1 as the current state-of-the-art, employing prompt designs informed by the OpenAI Cookbook for optimal few-shot performance. Detailed model architectures and prompt engineering strategies are discussed in Appendix˜D.
@@ 5.2 End-to-end Action Preference Learning
As shown in Table˜1 (the End-to-end row), video-based models produce Levenshtein distances approaching the average ground truth sequence lengths (15.80 at option level, 35.87 at sequence level). These high distances indicate that the models generate predominantly inconsistent action sequences, suggesting a failure to grasp preferences embedded in demonstration videos. While symbol-based models show modest improvements, their performance gains remain limited.
Table 1: Levenshtein distance between generated and ground truth action sequences. End-to-end represents models directly generating action sequences from demonstration-preference pair examples. Two-stage indicates generation using both demonstrations and previously inferred preference labels based on demonstrations. Second-stage (gt) uses demonstrations alongside ground truth preference labels for sequence generation.
@@
ViViT LLaVA-Next EILEV GPT-4V Llama3-8B DeepSeek-R1 GPT-4.1
Option Level
[…]
Sequence Level
[…]
Overall
End-to-end 24.76 25.35 22.99 24.69 23.26 18.72 16.95
Two-stage - 21.24 22.96 17.94 17.56 10.90 8.22
Second-stage (gt) - 11.10 18.88 6.31 13.62 7.93 5.23
@@ 5.3 Two-Stage Learning-Planning
Given the limitations of end-to-end learning, we implement a two-stage approach to decompose the preference learning problem. The first stage focuses on preference prediction, where we provide models with auxiliary preference token labels and train them to predict hidden preferences explicitly. These preference tokens, as discussed in Section˜4.1, maintain sufficient semantic content for translation into primitive actions.
Results from the first stage (Table˜2) reveal significant performance variations across models. At the option level, GPT-4V achieves superior performance with 48.48% accuracy, demonstrating strong capability in interpreting demonstrated preferences. Among symbol-based models, the stark contrast between DAG-Opt’s limited performance and the improved results from Llama3-8B and GPT-4.1 highlights the advantage of next-token prediction over dependency learning for preference inference. Models with language components consistently show improved preference understanding compared to end-to-end learning.
Table 2: Preference prediction accuracy in few-shot and ablative settings.
@@
ViViT LLaVA-Next EILEV GPT-4V DAG-Opt Llama3-8B DeepSeek-R1 GPT-4.1
Few-shot
Option Level 9.38 36.87 38.33 48.48 10.15 72.98 86.02 88.91
Sequence Level 4.24 24.85 32.69 37.50 13.49 67.18 71.21 70.28
Overall 6.81 30.86 35.51 42.99 11.82 70.08 78.62 79.60
Ablative
Option Level 9.16 15.47 4.77 29.42 3.84 39.50 78.19 75.29
Sequence Level 4.38 8.13 0.00 0.00 1.28 6.25 15.29 14.11
Overall 6.77 11.8 2.38 14.71 2.56 22.88 46.74 44.70
The second stage involves generating action sequences based on both demonstrations and predicted preference labels from the first stage, introducing potential error propagation. Results in Table˜1 (Two-stage row) and Figure˜5 (a)-(b) show significant improvements when models receive explicit preferences. For comprehensive evaluation, we include planning results using ground truth preference labels (Second-stage (gt) row). GPT-4V and GPT-4.1 achieve near-zero Levenshtein distances, indicating almost perfect alignment with ground truth action sequences.
@@
To isolate the impact of prior knowledge versus in-context learning, we conduct ablation studies by removing demonstrations and testing preference prediction on isolated test sequences. Results in Table˜2 (bottom) show significant performance degradation compared to few-shot learning (Table˜2 (top)), particularly at the sequence level. This suggests that while models may encode basic task-specific preferences, they rely heavily on demonstrations to recognize complex preference patterns in varied sequences.
@@
Figure 5: Levenshtein distance between generated and ground truth action sequences. Results shown for both (a) option level and (b) sequence level under two conditions: End-to-end bars represent direct sequence generation from previous observations, while Second-stage bars show performance when models receive predicted preference labels. The — line indicates average ground truth sequence length (option level: 15.80, sequence level: 35.87). Lower distances indicate better performance. Results demonstrate significantly improved performance under the two-stage approach compared to end-to-end generation.
@@ 5.4 Generalization
While human actions may vary across different objects and scenes, underlying preferences often remain consistent. We evaluate the models’ ability to generalize preference learning across varying visual contexts. The original test set inherently tests generalization by randomly sampling scenes and objects when rendering video demonstrations for each preference. To gain additional insights, we conduct complementary experiments with controlled conditions where demonstration and test videos are rendered in identical rooms with the same objects. This controlled setting enables direct performance comparisons under consistent conditions. We evaluate EILEV, LLaVA, and GPT-4.1 series models on this variant of PbP, as these models previously demonstrated strong few-shot reasoning capabilities. Results are summarized in Table˜3.
Table 3: Models’ generalization ability. direct denotes experiments conducted without generalization. gen denotes the original experiments conducted with generalization cases. Also the accuracy of preference prediction.
LLaVA-Next EILEV GPT-4V DeepSeek-R1 GPT-4.1
Option Level direct 33.25 46.93 53.24 86.02 88.91
Option Level gen 36.87 38.33 48.48 84.98 87.12
Sequence Level direct 33.12 37.53 39.42 71.21 70.28
Sequence Level gen 24.85 32.69 37.50 70.16 68.01
@@
Symbol-based reasoning (GPT-4.1) demonstrates consistent performance regardless of scene or object variations, while vision-based models show greater sensitivity to scene changes. This distinction stems from the nature of our predefined preferences, which are sufficiently abstract and general to apply across diverse scenes and objects. Vision-based models, however, tend to anchor their few-shot learned preferences to specific visual features of scenes or objects. When these visual elements change, preference recognition accuracy may deteriorate. This contextual dependence remains a persistent challenge for vision-based models, which often overfit to scene-specific features from training videos.
Analysis of test points across direct and gen conditions (Figure˜6) reveals two key findings: (i) Preference learning performance correlates with scene characteristics, with certain scenes proving consistently challenging across both conditions. (ii) While direct cases show better performance overall, failure patterns differ between conditions, particularly for vision-based models. This suggests models rely heavily on visual context consistency–including object arrangement and scene layout–for accurate predictions, indicating potential superficial learning rather than true preference understanding. Symbol-based reasoning maintains robust performance across varied scenes due to the general nature of predefined preferences, whereas vision-based models’ strong dependence on specific visual contexts limits their generalization capability.
@@ 5.5 Ablations on Demonstration Numbers
We examine the effect of demonstration quantity on model performance through an ablation study (Figure˜5 (c)-(d)). Results show that increasing demonstration numbers generally improves preference learning and planning effectiveness. This improvement is most evident in second-stage planning, where models achieve lower sequence distances by more accurately replicating human actions. Models like GPT-4.1, Llama3, and EILEV show consistent performance gains with additional demonstrations. However, we observe that excessive demonstrations (e.g., 5-demo cases for GPT-4.1 and EILVE) can sometimes impair first-stage prediction accuracy. Despite these occasional exceptions, the overall trend confirms our intuition: more demonstrations enhance learning and planning performance. These findings highlight the importance of demonstration quantity in developing effective personalized planning systems that align with user preferences.
@@ Appendix A Dataset Card
      Each instance contains an egocentric video of an agent’s activity, its bird’s-eye-view map of the position of the agent, and a frame-level textual annotation of the current action, as shown in Figure˜4. Additionally, we provide a rendered third-person view of the entire process.
@@ B.1 Preferenes
See Table˜A1 for the preference statistics in PbP.
@@ B.2 Actions
See Table˜A2 for the action statistics in PbP. We implement 17 action primitives in PbP to assist with model planning and dataset rendering. These action primitives have parameters that simplify tasks and are considered the lowest-level actions. Each sub-task contains 8 to 20 such lowest-level actions. Generally, most of these actions consist of two parts: the robot movement part and the arm (gripper) execution part. For robot movement, we use the A* algorithm to find paths and avoid collisions. We build a connection map during scene initialization for navigation, taking the robot’s width into consideration. For the arm (gripper) execution, we primarily use the IK algorithm to compute arm movements. However, since IK cannot handle complex tasks, such as picking objects from the fridge, we also leverage the Open Motion Planning Library (OMPL) planner (Sucan et al.,, 2012) with forward planning to assist in planning the arm positions.
@@ B.3 More Dataset Details and Discussion
The process of producing data is mainly explained in Section˜4.2. In summary, we follow the order of “sample preference - sample scene - sample objects to be manipulated - generate actions guided by a rule-based planner”.
@@
Considering that preferences in household activities are not only multi-dimensional but also hierarchical, we first define a hierarchy of preferences from the perspective of how things happen in a life scenario, that is, from each specific action to a sub-task consisting of several actions, and then to the sequence combining these sub-tasks. The next step is to expand each level with typical tasks and actions. The detailed definition of the 290 preferences can be found in Section˜4.1.
@@ D.6 Llama3-8B
D.7 GPT-4.1
We use GPT-4.1 from OpenAI with a temperature of 0.05. The textual prompts are as follows: